OpenStudy (anonymous):

Evaluate the indefinite integral. (3x)/(x^2+4)

OpenStudy (anonymous):

Let \(u=x^2+4\). What is \(dx\) equal to?

OpenStudy (anonymous):

dx= du/ (x^2+4)?

OpenStudy (anonymous):

\[u=x^2+4\] \[du=2xdx\] \[dx=\frac{du}{2x}\]

OpenStudy (anonymous):

oh yeah, i totally forgot that you have to take the derivative first. kay what's next?

OpenStudy (anonymous):

\[u=x^2+4\] \[\frac{du}{dx}=2x\] \[du=2xdx\] is another way to look at it

OpenStudy (anonymous):

plug that into the equation now

OpenStudy (anonymous):

\[\int \frac{3x}{x^2+4}\frac{du}{2x}\]

OpenStudy (anonymous):

but \[x^2+4=u\] \[\int\frac{3x}{u}\frac{du}{2x}\]

OpenStudy (anonymous):

now you shoudl be able to get something with u only and something smiple to solve

OpenStudy (anonymous):

so you take du/2x out right, cuz it's a constant or something like that?

OpenStudy (anonymous):

not quite you cant take variables out , but if you look the x's cancel and you're left with \[\int \frac{3}{2}\frac{1}{u}du\]

OpenStudy (anonymous):

constants you can pull out so you're left with a pretty simple integral \[\int \frac{du}{u}\]

OpenStudy (anonymous):

ok so then what do you plug in?

OpenStudy (anonymous):

what do you mean? indefinite integrals don't have upper nad lower limits

OpenStudy (anonymous):

like don't you have to integrate u? then plug it in?

OpenStudy (anonymous):

yes so you integrated it correct?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

um x^3/3+4x?

OpenStudy (anonymous):

or u^2/2?

OpenStudy (anonymous):

and what happens to du?

OpenStudy (agent0smith):

Do you know the integral of 1/x? \[\int\limits_{}^{} \frac{ 1 }{ x}dx = \ln \left| x \right|\] Here we have \[\int\limits \frac{1}{u} du = ?\]

OpenStudy (anonymous):

yeah...

OpenStudy (agent0smith):

So then what's the integral of \[\int\limits\limits \frac{1}{u} du = ? \]

OpenStudy (agent0smith):

Given you know that \[\int\limits\limits_{}^{} \frac{ 1 }{ x}dx = \ln \left| x \right| \]

OpenStudy (agent0smith):

Oh don't forget the 3/2 out front...\[\frac{3}{2} \int\limits \frac{1}{u}du \]

OpenStudy (anonymous):

I don't think you comprehend the point of a substitution.